# Liquidity fragmentation on decentralized exchanges

### v5

@@ 5 Conclusion
We show, both theoretically and empirically, that fixed costs of liquidity management promote market fragmentation across decentralized pools and generate clienteles of liquidity providers. Large market makers, likely institutions and funds, have stronger economies of scale and can afford to frequently manage their positions on very active low fee markets. On the other hand, smaller retail liquidity providers become confined to high fee markets with scant activity, trading off a lower execution probability against lower gas costs to update their positions. Since large liquidity providers can churn their position at a faster pace, two thirds of the trading volume interacts with less than half the capital locked on Uniswap V3.
@@ Appendix B Notation summary
$v$, $\Delta$ Common and private values of the token.
@@
$q_{i}$ Token endowment of liquidity provider $i$.
$\varphi\left(q,Q\right)$ Distribution density of LP endowment $q$.
$Q$ Maximum token endowment for LPs.
@@
$\theta$ Linear trading rate of small LTs.
$\lambda$ Poisson arrival rate of large LTs.
@@
$\mathcal{L}_{k}$ Equilibrium liquidity supply on exchange $k\in\left\{\textbf{L},\textbf{H}\right\}$.
@@
$\underline{q}$ Lowest token endowment deposited on the market (from break-even condition).
$d_{k}$ Duration of a liquidity cycle on exchange $k$.
@@ Appendix C Proofs
We rewrite equation (2.1) as
$\pi_{L}-\pi_{H}$ $=\frac{1}{\lambda d_{H}d_{L}}\left[q_{i}\left(h\times\left(\frac{q_{\text{t}}}{Q}\right)^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}-\left(h-l\right)\right)-\Gamma\left(\frac{q_{\text{t}}}{Q}\right)^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\right].$ (C.1)
If the coefficient of $q_{i}$ in (C.1) is negative, that is $h\times\left(\frac{q_{\text{t}}}{Q}\right)^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}-\left(h-l\right)\leq 0$, then $\pi_{L}-\pi_{H}<0$ and all LPs optimally choose to post liquidity on pool $H$. A necessary condition for a fragmented equilibrium is therefore that
$f_{1}\left(q_{\text{t}}\right)\coloneqq h\times\left(\frac{q_{\text{t}}}{Q}\right)^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}-\left(h-\ell\right)>0$ (C.2)
The function $f_{1}\left(q_{\text{t}}\right)$ increases in $q_{\text{t}}$, and therefore it has at most one root $q_{r}$ in $\left[1,Q\right]$. Since $f_{1}\left(Q\right)=\ell>0$, it follows that $q_{r}\in\left[1,Q\right)$ if and only if
$f_{1}\left(1\right)<0\Rightarrow\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}>1$ (C.3)
Second, in any fragmented market equilibrium, the marginal LP is indifferent between providing liquidity on the $L$ or $H$ pool. Equivalently, the marginal trader solves $\pi_{L}-\pi_{H}=0$, which from equation (11) translates to a solution of
$f_{2}\left(q_{\text{t}}\right)\coloneqq\Gamma\frac{q_{\text{t}}^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\times Q^{-\frac{\lambda}{\theta}\frac{Q}{Q-1}}}{h\left[q_{\text{t}}^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\times Q^{-\frac{\lambda}{\theta}\frac{Q}{Q-1}}\right]-\left(h-\ell\right)}-q_{\text{t}}=0.$ (C.4)
We show that $f_{2}\left(\cdot\right)$ monotonically decreases in $q_{\text{t}}$ since
$\frac{\partial f_{2}\left(q_{\text{t}}\right)}{\partial q_{\text{t}}}=-\Gamma\frac{\lambda(h-l)Q^{\frac{\lambda Q}{\theta(Q-1)}+1}q_{\text{t}}^{\frac{\lambda Q}{\theta(Q-1)}-1}}{\theta(Q-1)\left((l-h)Q^{\frac{\lambda Q}{\theta(Q-1)}}+hq_{\text{t}}^{\frac{\lambda Q}{\theta(Q-1)}}\right)^{2}}-1<0,$ (C.5)
and therefore it has at most one root. That is, if a fragmented market equilibrium exists, then it is unique.
To complete the proof, we consider two cases. First, let $\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}>1$ such that $q_{r}>1$. For a belief $q_{\text{t}}<q_{r}<Q$, liquidity providers optimally choose pool $H$ – which is equivalent to having an outcome in which the marginal trader is $q_{\text{t}}=Q$. Therefore, there is no fulfilled expectation equilibrium on $\left[1,q_{r}\right]$ in which beliefs match outcomes. From the intermediate value theorem and monotonicity of $f_{2}\left(q_{\text{t}}\right)$, there is exactly one root of $f_{2}\left(q_{\text{t}}\right)$ on $\left(q_{r},Q\right]$ since
$\underbrace{\frac{\Gamma}{\ell}-Q}_{<0}=f_{2}\left(Q\right)<f_{2}\left(q_{\text{t}}\right)<\lim_{q_{\text{t}}\downarrow q_{r}}f_{2}\left(q_{\text{t}}\right)=\infty.$ (C.6)
Therefore, for $\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}>1$ there is one unique equilibrium with fragmented markets.
Second, consider the case where $\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\leq 1$ such that $q_{r}<1$. In this case the reaction function $f_{2}\left(q_{\text{t}}\right)$ is decreasing and has no discontinuities. If $\frac{\Gamma}{h}>1$ such that the marginal LP earns zero expected profit, it follows from the intermediate value theorem and monotonicity of $f_{2}$ that there is a unique fragmented equilibrium since $f_{2}\left(\frac{\Gamma}{h}\right)>0$ and $f_{2}\left(Q\right)<0$.
Otherwise – if $\frac{\Gamma}{h}<1$ – a necessary and sufficient condition for a unique interior root is that $f_{2}\left(1\right)>0$ or, equivalently, that
$\frac{\Gamma}{h}>1-\frac{h-l}{h}Q^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}.$ (C.7)
If (C.7) does not hold, then $f_{2}\left(1\right)<0$ which, from monotonicity, implies that $f_{2}\left(q_{\text{t}}\right)<0$ for all $q_{\text{t}}\in\left[1,Q\right]$. In this case, the unique equilibrium is that all liquidity providers deposit their endowments on pool $L$. ∎
Lemma C.1.
[…]
Lemma C.1
Proof.
From the implicit function theorem,
$\frac{\partial q_{\text{t}}^{\star}}{\partial x}=-\left(\frac{\partial f_{2}}{\partial q_{\text{t}}}\right)^{-1}\frac{\partial f_{2}}{\partial x}.$ (C.8)
Since from equation (C.5) $f_{2}\left(\cdot\right)$ decreases in $q_{\text{t}}$, it follows that the sign of $\frac{\partial q_{\text{t}}^{\star}}{\partial x}$ is the same as the sign of $\frac{\partial f_{2}}{\partial x}$, which significantly simplifies the task.
First, we note that
$\frac{\partial f_{2}}{\partial\Gamma}=\frac{q_{\text{t}}^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\times Q^{-\frac{\lambda}{\theta}\frac{Q}{Q-1}}}{h\left[q_{\text{t}}^{\frac{\lambda}{\theta}\frac{Q}{Q-1}}\times Q^{-\frac{\lambda}{\theta}\frac{Q}{Q-1}}\right]-\left(h-\ell\right)}>0,$ (C.9)
and therefore $q_{\text{t}}^{\star}$ increases in $\Gamma$. Second,
$\frac{\partial f_{2}}{\partial\lambda}=\frac{\Gamma(h-l)Q^{\frac{\lambda Q}{\theta(Q-1)}+1}(\log(Q)-\log(q_{\text{t}}))q_{\text{t}}^{\frac{\lambda Q}{\theta(Q-1)}}}{\theta(Q-1)\left((l-h)Q^{\frac{\lambda Q}{\theta(Q-1)}}+hq_{\text{t}}^{\frac{\lambda Q}{\theta(Q-1)}}\right)^{2}}>0,$ (C.10)
since $q_{\text{t}}<Q$. Therefore, $q_{\text{t}}^{\star}$ increases in $\lambda$. Third,
$\frac{\partial f_{2}}{\partial\theta}=-\frac{\Gamma\lambda(h-l)Q^{\frac{\lambda Q}{\theta(Q-1)}+1}(\log(Q)-\log(q_{\text{t}}))q_{\text{t}}^{\frac{\lambda Q}{\theta(Q-1)}}}{(Q-1)\left(\theta(h-l)Q^{\frac{\lambda Q}{\theta(Q-1)}}-h\theta q_{\text{t}}^{\frac{\lambda Q}{\theta(Q-1)}}\right)^{2}}<0,$ (C.11)
which implies that $q_{\text{t}}^{\star}$ decreases in $\theta$.
Finally, we compute partial derivatives with respect to pool fees $h$ and $l$. That is
$\frac{\partial f_{2}}{\partial h}=\frac{\Gamma Q^{-\frac{\lambda Q}{(Q-1)\theta}}\left(q_{\text{t}}^{-\frac{\lambda Q}{(Q-1)\theta}}-Q^{-\frac{\lambda Q}{(Q-1)\theta}}\right)}{\left((l-h)q_{\text{t}}^{\frac{\lambda Q}{\theta-\theta Q}}+hQ^{\frac{\lambda Q}{\theta-\theta Q}}\right)^{2}}>0$ (C.12)
since $q_{\text{t}}<Q$ and $\frac{\lambda Q}{(Q-1)\theta}>0$; further,
$\frac{\partial f_{2}}{\partial l}=-\frac{\Gamma Q^{\frac{\lambda Q}{\theta-\theta Q}}\text{qm}^{\frac{\lambda Q}{\theta(Q-1)}}}{\left(h\left(Q^{\frac{\lambda Q}{\theta-\theta Q}}\text{qm}^{\frac{\lambda Q}{\theta(Q-1)}}-1\right)+l\right)^{2}}<0.$ (C.13)
@@
The comparative statics with respect to $\lambda$, $\theta$, and $\ell$ follow immediately from Lemma C.1 and the fact that $w_{L}$ decreases in $q_{\text{t}}$.
Comparative statics for $h$.
If $\frac{\Gamma}{h}>1$ such that $\underline{q}=\frac{\Gamma}{h}$, then the partial derivative of $w_{L}$ with respect to $h$ is
$\frac{\partial w_{L}}{\partial h}=-\frac{\left(\log Q-\log q_{\text{t}}\right)q_{\text{t}}+h\left(\log q_{\text{t}}-\log\frac{\Gamma}{h}\right)\frac{\partial q_{\text{t}}}{\partial h}}{h\left(\log Q-\log q_{\text{t}}\right)^{2}q_{\text{t}}},$ (C.14)
which is positive since from Lemma C.1 we have that $\frac{\partial q_{\text{t}}}{\partial h}>0$. Otherwise, if $\frac{\Gamma}{h}>1$ such that $\underline{q}=1$, then
$\frac{\partial w_{L}}{\partial h}=-\frac{\partial q_{\text{t}}}{\partial h}\frac{1}{\log Qq_{\text{t}}}<0.$ (C.15)
Comparative statics for $\Gamma$.
If $\underline{q}=\frac{\Gamma}{h}$ (i.e., if $\frac{\Gamma}{h}>1$), then the partial derivative of $w_{L}$ with respect to $\Gamma$ is
$\frac{\partial w_{L}}{\partial\Gamma}=\frac{\Gamma\frac{\partial q_{\text{t}}}{\partial\Gamma}\left(\log\left(\frac{\Gamma}{h}\right)-\log(Q)\right)+q_{\text{t}}(\log(Q)-\log(q_{\text{t}})))}{\Gamma q_{\text{t}}\left(\log(Q)-\log\left(\frac{\Gamma}{h}\right)\right)^{2}}.$ (C.16)
From equation (11), $q_{\text{t}}$ is linear in $\Gamma$ and therefore $q_{\text{t}}=\Gamma\frac{\partial q_{\text{t}}}{\partial\Gamma}$. We can rewrite (C.16) as
$\frac{\partial w_{L}}{\partial\Gamma}=-\frac{\log\left(q_{\text{t}}\right)-\log\left(\frac{\Gamma}{h}\right)}{\Gamma\left(\log\left(Q\right)-\log\left(\frac{\Gamma}{h}\right)\right)^{2}}<0.$ (C.17)
Further, if $\underline{q}=1$, then it is easy to see that $w_{L}$ decreases in $\Gamma$ since $q_{\text{t}}$ increases with $\Gamma$ and the denominator is not a function of $\Gamma$.
@@
The two-pool implementation shortfall is
$\text{IS}\left(\left\{h,\ell\right\}\right)=\ell\int_{q_{\text{t}}}^{Q}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i+h\int_{\frac{\Gamma}{h}}^{q_{\text{t}}}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i+g\left(\Theta-\int_{\frac{\Gamma}{h}}^{Q}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i\right).$ (C.18)
We set $h=f$ such that the marginal LP entering the market is the same as in the single-fee pool, which equates the opportunity cost IS component across the two setups. It follows that:
$\text{IS}\left(\left\{f,\ell\right\}\right)-\text{IS}\left(\left\{f\right\}\right)=\left(\ell-f\right)\int_{q_{\text{t}}}^{Q}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i\leq 0,$ (C.19)
with strict inequality if $q_{\text{t}}<Q$ such that the low-fee pool attracts a positive mass of LPs.∎

### v7

@@ 5 Conclusion
We show, both theoretically and empirically, that fixed costs of liquidity management promote market fragmentation across decentralized pools and generate clienteles of liquidity providers. Large market makers, likely institutions and funds, have stronger economies of scale and can afford to frequently manage their positions on very active low fee markets, while bearing higher adverse selection risk. On the other hand, smaller retail liquidity providers become confined to high fee markets with scant activity, trading off a lower execution probability against reduced adverse selection and lower gas costs to update their positions. Since large liquidity providers can churn their position at a faster pace, two thirds of the trading volume interacts with less than half the capital locked on Uniswap V3.
@@ Appendix B Notation summary
$v_{t}$ Common value of the token at time $t$.
$\eta$, $1-\eta$ Poisson arrival rate of news and private value shocks, respectively.
$\delta$ Size of common or private value shock.
$\Delta$ Parameter governing the probability distribution of shocks, $\phi\left(\delta\right)=\frac{1}{2\Delta\sqrt{1+\delta}}$.
@@
$q_{i}$ Token endowment of liquidity provider $i$, exponentially distributed with scale parameter $\lambda$.
$\lambda$ Aggregate liquidity supply if all LPs join the market.
@@
$r$ Width of the price range $\left[\frac{v}{\left(1+r\right)^{2}},v(1+r)^{2}\right]$.
@@
$T_{k}$ Equilibrium liquidity supply on exchange $k\in\left\{\textbf{L},\textbf{H}\right\}$.
$\tau\left(\delta\right)$ Optimal trade size for LT or A with value shock $\delta$.
$\mathcal{L}\left(f_{k}\right)$ Liquidity yield (fee revenue per unit of liquidity supplied) on pool with fee $f_{k}$.
$\mathcal{A}\left(f_{k}\right)$ Adverse selection cost for LP per unit of liquidity supplied on pool with fee $f_{k}$.
$\mathcal{C}\left(f_{k}\right)$ Liquidity re-balancing cost on pool with fee $f_{k}$.
@@
$\underline{q}_{k}$ Lowest token endowment deposited on pool $k$ (from break-even condition).
@@ Appendix C Proofs
Lemma 1
Proof.
We take the expectation of fee revenues over the size of private value shocks $\delta$ and obtain:
[…]
$\mathbb{P}\left(f_{k}<\delta\leq(1+f_{k})(1+r)^{2}-1\right)\times\frac{1+r}{r}\mathbb{E}\left[\sqrt{\frac{1+\delta}{1+f_{k}}}-1\mid f_{k}<\delta\leq(1+f_{k})(1+r)^{2}-1\right]+$
$+\mathbb{P}\left(\delta>(1+f_{k})(1+r)^{2}-1\right)\times\left(1+r\right)\Big{\}}$
$=q_{i}\underbrace{v\frac{f_{k}(r+1)\left(2\Delta-r\sqrt{f_{k}+1}-2\sqrt{f_{k}+1}\right)}{\Delta}}_{\equiv\mathcal{L}\left(f_{k}\right)},$ (C.1)
where we define $\mathcal{L}\left(f_{k}\right)$ as the liquidity yield: i.e., the per-unit profit from liquidity provision in pool $k$.
To explore how the liquidity revenue changes with respect to the pool fee, we differentiate $\mathcal{L}$ with respect to $f$:
$\frac{\partial{L}\left(f\right)}{\partial f}=-\frac{(r+1)\left(2\left(-2\Delta\sqrt{f+1}+r+2\right)+3f(r+2)\right)}{4\Delta\sqrt{f+1}}.$ (C.2)
Starting with $f=0$ and given that $\Delta>1+r$ (by Assumption 1), the derivative at $f=0$ is positive:
$\frac{\partial\mathcal{L}(f)}{\partial f}\bigg{|}_{f=0}=\frac{(r+1)(2\Delta-r-2)}{2\Delta}>0,$ (C.3)
indicating that liquidity revenue increases with pool fee at this point. The derivative has roots:
$f_{1,2}=\frac{-6(r+2)^{2}+8\Delta^{2}\pm 4\sqrt{4\Delta^{4}+3\Delta^{2}(r+2)^{2}}}{9(r+2)^{2}},$ (C.4)
where the smallest root $f_{1}$ is negative and therefore not relevant. We need to show that the largest root $f_{2}$ is always positive, defining the threshold $\overline{f}$.
For this, consider the numerator of $f_{2}$, labeled $g(r,\Delta)$:
$g(r,\Delta)=8\Delta^{2}+4\sqrt{4\Delta^{4}+3\Delta^{2}(r+2)^{2}}-6(r+2)^{2}.$
This function has three roots in $r$, all of which are negative: $r=-2$, $r=-2(1+\Delta)$, and $r=2(\Delta-1)$. Since these roots are negative, for $r\geq 0$, $g$ does not change sign and it is sufficient to examine $g(0,\Delta)$:
$g(0,\Delta)=8\Delta^{2}+4\sqrt{\Delta^{2}(\Delta^{2}+3)}-6.$ (C.5)
This is positive for any $\Delta\geq 1$, confirming that the largest root $f_{2}$ is positive and hence, $\overline{f}$ exists and is positive. This completes the proof that the liquidity revenue increases with the pool fee until $\overline{f}$ and decreases with pool fee for $f>\overline{f}$. ∎
Lemma 2
Proof.
The cost of adverse selection for pool $k$ after evaluating equation (2.1.3) is
$\mathcal{A}\left(f_{k}\right)=v\frac{\left(\Delta-\sqrt{1+f}(1+r)\right)\left(\Delta^{2}+\Delta\sqrt{f+1}\left(1+r\right)+(f+1)(r-2)(r+1)\right)+(f+1)^{3/2}r^{2}(r+1)}{3\Delta}.$ (C.6)
We aim to demonstrate that $\mathcal{A}(f)$ decreases as $f$ increases. To do this, we calculate the partial derivative of $\mathcal{A}(f)$ with respect to $f$:
$\frac{\partial\mathcal{A}\left(f\right)}{\partial f}=\frac{(r+1)\left(-2\Delta\sqrt{f+1}+f(r+2)+r+2\right)}{2\Delta\sqrt{f+1}}<0$ (C.7)
The derivative is negative if $\Delta>\frac{1}{2}\sqrt{1+f}(2+r)$. Given that $\frac{1}{2}(2+r)<1+r$, it follows that $\mathcal{A}(f)$ decreases for any $\Delta>(1+r)\sqrt{1+f}$, consistent with our assumption on $\Delta$. ∎
@@
First, consider the case in which $\eta>\frac{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)}{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)+\mathcal{A}\left(l\right)-\mathcal{A}\left(h\right)}$ holds. This implies $q_{\text{t}}<0$, and consequently, $\pi_{L}-\pi_{H}<0$ for all $q$. Under this scenario, liquidity providers universally favor pool $H$ over pool $L$. They supply liquidity on pool $H$ if and only if their participation constraint is satisfied, that is if $q_{i}>\underline{q}_{h}$.
Conversely, if $\eta\leq\frac{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)}{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)+\mathcal{A}\left(l\right)-\mathcal{A}\left(h\right)}$, then $q_{\text{t}}\geq 0$, allowing for a fragmented equilibrium. If $q_{\text{t}}\geq 0$, then $\pi_{L}$ has a steeper slope compared to $\pi_{H}$: profit increases more rapidly with liquidity supply in pool $L$ than in pool $H$. There are two potential outcomes.
  1. 1.
    Dominance of pool $L$. If $0<q_{\text{t}}<\underline{q}_{\ell}<\underline{q}_{h}$, as shown in the left-hand side panel of the diagram below, then the low-fee pool $L$ captures the entire market share for any $q_{i}$ that yields positive profits. The condition $\underline{q}_{\ell}<\underline{q}_{h}$ is equivalent to
    $\underline{q}_{\ell}<\underline{q}_{h}\Leftrightarrow\frac{\mathcal{C}\left(h\right)}{\mathcal{C}\left(\ell\right)}>\frac{\left(1-\eta\right)\mathcal{L}\left(h\right)-\eta\mathcal{A}\left(h\right)}{\left(1-\eta\right)\mathcal{L}\left(\ell\right)-\eta\mathcal{A}\left(\ell\right)},$ (C.8)
    which translates to $\eta>\frac{\mathcal{C}\left(\ell\right)\mathcal{L}\left(h\right)-\mathcal{C}\left(h\right)\mathcal{L}\left(\ell\right)}{\mathcal{C}\left(\ell\right)\left[\mathcal{L}\left(h\right)+\mathcal{A}\left(h\right)\right]-\mathcal{C}\left(h\right)\left[\mathcal{L}\left(\ell\right)+\mathcal{A}\left(\ell\right)\right]}$. Since we require $\eta\leq\frac{\mathcal{L}\left(\ell\right)}{\mathcal{L}\left(\ell\right)+\mathcal{A}\left(\ell\right)}$ by Assumption 2, it must be that $\mathcal{L}\left(\ell\right)\mathcal{A}\left(h\right)<\mathcal{L}\left(h\right)\mathcal{A}\left(\ell\right)$.
    However, the parameter regions never overlap, ruling out this scenarios: we will show that $\mathcal{L}\left(\ell\right)\mathcal{A}\left(h\right)-\mathcal{L}\left(h\right)\mathcal{A}\left(\ell\right)>0$. To see this, we first note that $\eta\leq\frac{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)}{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)+\mathcal{A}\left(l\right)-\mathcal{A}\left(h\right)}$ is equivalent to:
    $\eta\leq\frac{3\left(h\left(2\Delta-\sqrt{h+1}r-2\sqrt{h+1}\right)-\ell\left(2\Delta-\sqrt{\ell+1}r-2\sqrt{\ell+1}\right)\right)}{(r+2)\left(\sqrt{h+1}\left(h-2\right)-\sqrt{\ell+1}\left(\ell-2\right)\right)},$ (C.9)
    which implies that $h\left(2\Delta-\sqrt{h+1}r-2\sqrt{h+1}\right)-\ell\left(2\Delta-\sqrt{\ell+1}r-2\sqrt{\ell+1}\right)>0$ since the denominator is always positive. We use the inequality in (C.9) and obtain that
    $\mathcal{L}\left(\ell\right)\mathcal{A}\left(h\right)-\mathcal{L}\left(h\right)\mathcal{A}\left(\ell\right)>\frac{1+r}{6\Delta^{2}}g,$ (C.10)
    where
    $g$ $\equiv\underbrace{(1+r)l\left(2\Delta-\sqrt{l+1}\left(r+2\right)\right)}_{>0}\times$ (C.11)
    ×$\times\left(\left(h-\ell\right)\Delta-(r+2)\left(\sqrt{h+1}-\sqrt{l+1}\right)+\underbrace{h\left(2\Delta-\sqrt{h+1}\left(r+2\right)\right)-l\left(2\Delta-\sqrt{l+1}\left(r+2\right)\right)}_{>0}\right)>0,$ (C.12)
    given Assumption 1 and equation (C.9). To see that $\left(h-\ell\right)\Delta-(r+2)\left(\sqrt{h+1}-\sqrt{l+1}\right)>0$, we first note the expression increases in $\Delta$ and is therefore larger than $\sqrt{h+1}(r+1)(h-l)+(r+2)\left(\sqrt{l+1}-\sqrt{h+1}\right)$ for $\Delta>\sqrt{1+h}\left(1+r\right)$. The latter expression increases in $h$ and equals zero for $h=\ell$.The latter expression increases in $h$ and equals zero for $h=\ell$.
    Therefore, there are no parameter values for which $0<q_{\text{t}}<\underline{q}_{\ell}<\underline{q}_{h}$ and $\eta\leq\frac{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)}{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)+\mathcal{A}\left(l\right)-\mathcal{A}\left(h\right)}$, which rules out the case of pool $L$ attracting full market share.
  2. 2.
    Fragmented Market Equilibrium. The right-hand side panel depicts the scenario $\underline{q}_{h}<\underline{q}_{\ell}<q_{\text{t}}$. Here, liquidity providers with $q_{i}$ in the range $(\underline{q}_{h},q_{\text{t}}]$ achieve higher positive profits in pool $H$, while those with $q_{i}>q_{\text{t}}$ obtain larger profits in pool $L$. The condition $\underline{q}_{h}<\underline{q}_{\ell}$ is equivalent to
    $\underline{q}_{h}\leq\underline{q}_{\ell}\Leftrightarrow\frac{\mathcal{C}\left(h\right)}{\mathcal{C}\left(\ell\right)}\leq\frac{\left(1-\eta\right)\mathcal{L}\left(h\right)-\eta\mathcal{A}\left(h\right)}{\left(1-\eta\right)\mathcal{L}\left(\ell\right)-\eta\mathcal{A}\left(\ell\right)},$ (C.13)
    which is always true if $\eta\leq\frac{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)}{\mathcal{L}\left(l\right)-\mathcal{L}\left(h\right)+\mathcal{A}\left(l\right)-\mathcal{A}\left(h\right)}$ as seen above.
$\pi_{k}$$\underline{q}_{\ell}$$\pi_{L}$$\mathcal{C}_{\ell}$$\underline{q}_{h}$$\pi_{H}$$\mathcal{C}_{h}$$q_{\text{t}}$(a) Single pool market (with fee $\ell$)$\pi_{k}$$\underline{q}_{\ell}$$\pi_{L}$$\mathcal{C}_{\ell}$$\underline{q}_{h}$$\pi_{H}$$\mathcal{C}_{h}$$q_{\text{t}}$(b) Fragmented market
It is crucial to note that configurations where $\underline{q}_{\ell}<q_{\text{t}}<\underline{q}_{h}$ or $\underline{q}_{h}<q_{\text{t}}<\underline{q}_{\ell}$ are not feasible, as they would lead to a contradiction where profits are simultaneously positive in one pool and negative in the other at the indifference point $q_{\text{t}}$.
@@
We first note that both $q_{\text{t}}$ and $\underline{q}_{h}$ scale linearly with $\Gamma$: that is, there exists $Q_{t}>Q_{h}>0$ such that $q_{\text{t}}=\Gamma Q_{t}$ and $\underline{q}_{h}=\Gamma Q_{h}$ where $Q_{t}$ and $Q_{h}$ are not functions of $\Gamma$. Next, we compute the partial derivative of $w_{\ell}$ with respect to $\Gamma$
$\frac{\partial w_{\ell}}{\partial\Gamma}=\frac{\Gamma(Q_{h}-Q_{t})e^{\frac{\Gamma(Q_{h}-Q_{t})}{\lambda}}(\Gamma Q_{h}Q_{t}+\lambda(Q_{h}+Q_{t}))}{\lambda(\lambda+\Gamma Q_{h})^{2}}<0,$ (C.14)
since $Q_{h}<Q_{t}$ and all other terms are positive.∎
@@
The two-pool gains from trade for an LT with private value $v\left(1+\delta\right)$ are
$\text{GainsFromTrade}\left(\left\{h,\ell\mid\delta\right\}\right)$ $=v\delta T_{H}\min\left\{1,\frac{1+r}{r}\max\left\{0,1-\sqrt{\frac{1+h}{1+\delta}}\right\}\right\}+$
$+v\delta T_{L}\min\left\{1,\frac{1+r}{r}\max\left\{0,1-\sqrt{\frac{1+\ell}{1+\delta}}\right\}\right\}.$ (C.15)
We set $h=f$ such that the marginal LP entering the market is the same as in the single-fee pool; that is, $T_{H}+T_{L}=e^{-q_{h}\lambda}\left(q_{h}+\lambda\right)$. Since $\min\left\{1,\frac{1+r}{r}\max\left\{0,1-\sqrt{\frac{1+f}{1+\delta}}\right\}\right\}$ decreases in $f$, it follows that:
$\text{GainsFromTrade}\left(\left\{h,\ell\mid\delta\right\}\right)$ $\geq v\delta\underbrace{\left(T_{H}+T_{L}\right)}_{=e^{-q_{h}\lambda}\left(q_{h}+\lambda\right)}\min\left\{1,\frac{1+r}{r}\max\left\{0,1-\sqrt{\frac{1+h}{1+\delta}}\right\}\right\}$ (C.16)
$=\text{GainsFromTrade}\left(\left\{h\right\}\mid\delta\right\},$ (C.17)
with strict inequality if $q_{\text{t}}<Q$ such that the low-fee pool attracts a positive mass of LPs. The inequality holds for any $\delta$, and therefore it remains true if we aggregate the gains from trade over the distribution LT private values. We note that the expected gains from trade per unit of liquidity is
$\mathbb{E}\min\left\{1,\frac{1+r}{r}\max\left\{0,1-\sqrt{\frac{1+h}{1+\delta}}\right\}\right\}-$
$=\frac{6\sqrt{f+1}(r+1)\log(r+1)+r\left(2\left(\Delta^{3}-3\Delta-\sqrt{f+1}\right)+\sqrt{f+1}(f(r+1)(r+2)+r(r+3))\right)}{6\Delta r}>0,$ (C.18)
which decreases in $f$ since
$\frac{\partial\mathbb{E}\min\left\{1,\frac{1+r}{r}\max\left\{0,1-\sqrt{\frac{1+h}{1+\delta}}\right\}\right\}}{\partial f}=-\frac{(r+1)((f+1)r(r+2)-2\log(r+1))}{4\Delta\sqrt{f+1}r}<0.$ (C.19)